# Variational Inference Optimized Using the Curved Geometry of Coupled Free Energy

### v1

@@ 5 Conclusion
We present a generalization of VI, using the Coupled Free Energy (negative of the Evidence Lower Bound), which enables the training of robust, complex models using robust samples. The method utilizes the geometric curvature induced by the heavy-tailed characteristics of the coupled exponential family of distributions. The example of the VAE is used to demonstrate the improvement in the ability to learn complex datasets. The robustness of the models is improved by using a heavy-tailed latent distribution (the coupled Gaussian) and by increasing the cost of outliers via the coupled Free Energy.
@@
An original element of the design is that the statistical robustness of the learning is improved by drawing samples from the coupled probability (also known as an escort probability) of the latent distribution. The coupled probability, which raises the distribution to the power $1+\frac{\alpha\kappa}{1+d\kappa}$ and renormalizes, reduces both the scale and shape of the distribution. Thus, although the process is learning a high-entropy distribution, the samples are being drawn from a significantly lower-entropy distribution. While consistent with the theory of nonextensive statistical mechanics, this is a novel mechanism in that the coupled entropy is a compromise between the Tsallis entropy, which doesn’t include the normalization, and the normalized Tsallis entropy, which is unstable.
@@
Preliminary evidence of the robustness of this CVAE design is provided by the significant reduction in the CFE with $\kappa=1$ versus the FE with $\kappa=0$ measurements. While these are different metrics given the change in $\kappa$, non-robust models would be susceptible to an increase over the FE metric. Additional measurements of the robustness are planned using corrupted images as input seeds. A 3% improvement in the accuracy of the reconstructions was measured using the Fréchet Inception Distance, which is equivalent to the Wasserstein-2 metric. Neither the VAE nor the CVAE was improved by training the off-diagonal elements of the correlation matrix. This may be do with the prior distribution having off-diagonal elements set to zero.
@@
Our future research will include computation of the curved gradients using the coupled affine connection to improve the speed and stability of the learning rate. This first demonstration utilized the standard Euclidean gradients of the coupled Free energy. Additional research objectives are to explore methods to determine criteria for an optimal coupling value and the limits of how high the coupling can be set. Negative coupling values can also be valuable in producing models intended to be more decisive and extreme. The coupled VAE is just one example of the many techniques in VI that can be generalized to undertake more complex AI tasks.

### v2

@@ 5 Conclusion
We present a generalization of VI, using the Coupled Free Energy (Coupled Evidence Lower Bound of the inverse probabilities), which enables the training of robust, complex models using robust samples. The method utilizes the geometric curvature induced by the heavy-tailed characteristics of the coupled exponential family of distributions. The example of the VAE is used to demonstrate the improvement in the ability to learn complex datasets. The robustness of the models is improved by using a heavy-tailed latent distribution (the coupled Gaussian) and by increasing the cost of outliers via the coupled Free Energy.
@@
An original element of the design is that the statistical robustness of the learning is improved by drawing samples from the coupled probability (also known as an escort probability) of the latent distribution. The coupled probability, which raises the distribution to the power $q=1+\frac{\alpha\kappa}{1+d\kappa}$ and renormalizes, reduces the distribution’s scale and shape. The coupled probability samples are equivalent to conditioning the original distribution to subselect for samples in which $q$ of them are equal. The effect is to filter extreme outliers from the training process, while still assuring that the learned model will be robust against those outliers. While consistent with the theory of nonextensive statistical mechanics, this is a novel mechanism in that the coupled entropy is a compromise between the Tsallis entropy, which doesn’t include the normalization, and the normalized Tsallis entropy, which is unstable.
@@
Preliminary evidence of the robustness of this CVAE design is provided by the significant reduction in the CFE with $\kappa=1$ versus the FE with $\kappa=0$ measurements. While these are different metrics given the change in $\kappa$, non-robust models would be susceptible to an increase over the FE metric. Additional measurements of the robustness are planned using corrupted images as input seeds. A 3% improvement in the accuracy of the reconstructions was measured using the Fréchet Inception Distance, which is equivalent to the Wasserstein-2 metric. Neither the VAE nor the CVAE was improved by training the off-diagonal elements of the correlation matrix. This may be due to the prior distribution having off-diagonal elements set to zero.
While the Coupled Free Energy function derivation is complex, the machine learning implementation is relatively simple for a couple of reasons.
  • First, the coupled Gaussian, which is equivalent to the Student’s t, is a well-established heavy-tailed model. It was first used by Guinness Brewing Company to improve their beer production in the early 1900s.
  • Second, we establish how to train quite extreme heavy-tailed models (coupling = 1 is a Cauchy distribution, which has no well-defined moments), while drawing samples from significantly faster decaying distributions (1/3 in the Cauchy case, which has a well-defined first and second moment).
  • The reconstruction likelihood is proven to be the same as the original Free Energy, a mean-square average, with just a modification of the constants. It’s even possible that these constants are not a necessary part of the training, though we haven’t validated this yet.
@@
Our future research will include the computation of the curved gradients using the coupled affine connection to improve the speed and stability of the learning rate. This first demonstration utilized the standard Euclidean gradients of the coupled Free energy. Additional research objectives are to explore methods to determine criteria for an optimal coupling value and the limits of how high the coupling can be set. Negative coupling values can also be valuable in producing models intended to be more decisive and extreme. The coupled VAE is just one example of the many techniques in VI that can be generalized to undertake more complex AI tasks.